# Towards Real-World Stickers Use: A New Dataset for Multi-Tag Sticker Recognition

### v2

@@ Abstract
In real-world conversations, the diversity and ambiguity of stickers often lead to varied interpretations based on the context, necessitating the requirement for comprehensively understanding stickers and supporting multi-tagging. To address this challenge, we introduce StickerTAG, the first multi-tag sticker dataset comprising a collected tag set with 461 tags and 13,571 sticker-tag pairs111We believe that the release of this dataset will provide exciting research opportunities and encourage further research in sticker analysis. , designed to provide a deeper understanding of stickers. Recognizing multiple tags for stickers becomes particularly challenging due to sticker tags usually are fine-grained attribute aware. Hence, we propose an Attentive Attribute-oriented Prompt Learning method, i.e., AttPL, to capture informative features of stickers in a fine-grained manner to better differentiate tags. Specifically, we first apply an Attribute-oriented Description Generation (ADG) module to obtain the description for stickers from four attributes. Then, a Local Re-attention (LoR) module is designed to perceive the importance of local information. Finally, we use prompt learning to guide the recognition process and adopt confidence penalty optimization to penalize the confident output distribution. Extensive experiments show that our method achieves encouraging results for all commonly used metrics.
@@
Bingbing Wang, Bin Liang, Chun-Mei Feng, Wangmeng Zuo, Zhixin Bai, Shijue Huang, Kam-Fai Wong, Ruifeng Xu††thanks: Corresponding Authors Harbin Institute of Technology, Shenzhen, China Harbin Institute of Technology, Harbin, China The Chinese University of Hong Kong, Hong Kong, China Institute of High Performance Computing, Singapore {bingbing.wang, bin.liang, 2021113309}@stu.hit.edu.cn, kfwong@se.cuhk.edu.hk, {wmzuo,xuruifeng}@hit.edu.cn, {strawberry.feng0304,joehsj310}@gmail.com
@@ 1 Introduction
To tackle this challenge, we design an Attentive Attribute-oriented Prompt Learning method, abbreviated as AttPL, which begins with an Attribute-oriented Description Generation (ADG) module powered by a Multi-modal Large Language Model (MLLM) to derive the description of sticker based on the content, style, role, and action. Furthermore, the sticker is fed into a Local Re-attention (LoR) module for deriving patch-attentive embedding. Then, the recognition process is guided by prompt classification, where we initialize the soft prompt with attribute-oriented descriptions. Finally, we implement a confidence penalty optimization for precise and reliable recognition. We employ commonly used metrics to evaluate the performance of our model, and extensive experiments conducted on our StickerTAG dataset show that our approach significantly outperforms the strong baselines. Our contributions are summarized as follows:
@@
Extensive experiments conducted on the StickerTAG dataset show that our model outperforms all baselines, clearly confirming its effectiveness.
@@ 2.2 Sticker-based Method
Unlike image-text pairs in vision-language datasets, stickers demand finer-grained emotion recognition and subject understanding due to their increased diversity and domain specificity. This has prompted the development of specialized approaches tailored to the unique attributes of stickers, as relying on generic vision-language models (VLMs) may prove suboptimal. CSMSA (Ge et al., 2022) introduces a Sticker-Aware Multi-modal Sentiment Analysis Model to tackle challenges in multimodality, inter-series variations, and multi-modal sentiment fusion. Zhao et al. (2023) delves into aligning visual and textual features with emotional cues and artificial painting. It investigates the feasibility of seamlessly integrating additional tools into Large Language Models (LLMs). However, simple classification methods are insufficient to address the challenges posed by multi-tag sticker recognition. This prompts us to delve into more fine-grained features of stickers and enhances the ability of the model to differentiate between tags more effectively.
@@ 3 StickerTAG Dataset
In this section, we provide a detailed description of our dataset construction process. Firstly, we discuss the process of tag construction, illustrated in Figure 2 in Section 3.1. Next, we outline the procedure for collecting and annotating stickers in Section 3.2. Finally, we analyze the characteristics of our dataset in Section 3.3.
@@ 3.3 Characteristics
The final dataset contains 461 tags and 13,571 sticker-tag pairs annotated data, and the large-scale dataset is for future work. This dataset is the first annotated sticker dataset for multi-tag recognition to the best of our knowledge. We present the word cloud distribution of sticker tags in Figure 3 (a). We can conclude that certain tags prominently feature emotionally charged and conversational words, reflecting a diverse range of topics. Moreover, Figure 3 (b) shows the number samples per tag. It can be seen that our dataset exhibits a long-tailed distribution like most of the multi-label datasets. We also conduct studies on the tags of each sticker. According to our statistics, nearly 0.35% stickers have 6 tags, 2.67% of 5 tags, 9.23% of 4 tags, 14.99 % of 3 tags, 25.31% of 2 tags, and 47.45% of a tag. And the average tag length is 3.10 words.
@@ 4 Method
This section begins with a concise introduction to the multi-tag sticker recognition task, followed by a detailed explanation of our proposed AttPL framework. As demonstrated in Figure 4, AttPL consists primarily of four components: 1) Attribute-oriented Description Generation designs prompts based on the attributes of the stickers for MLLM to generate attribute-oriented descriptions. 2) Local Re-attention Module utilizes masked images in a novel manner to derive renewed attention for deriving patching-attentive embedding. 3) Prompt-based Classification exploits attribute-oriented descriptions to initialize the soft prompts and combines patch-attentive embedding to achieve the multi-tag prediction. 4) Confidence Penalty Optimization penalizes confident output distributions for implicit feedback of renewed attention.
@@ 4.5 Confidence Penalty Optimization
To enhance the ability of LoR to capture important local information, we deploy a confidence penalty optimization method. Specifically, we first obtain the original image representation by passing the image patch embedding through an image encoder. Then, concatenated with soft prompts initialized by four description representations to derive the original image representation , which is fed into the classifier, yielding the original confidence distribution. We compare the probability changes in reconstructed confidence and original confidence for the correct tag. We calculate the sum of the magnitudes of the decreases in reconstructed confidence relative to original confidence and incorporate it as a penalty into the loss.
@@
5 Experiment
5.1 Experimental Set-up
Implement details. For data participation, we divide our StickerTAG dataset into a training set, a validation set, and a testing set with a ratio of 8:1:1. AttPL employs Swin Transformer Vaswani et al. (2017) as the image encoder, we apply a convolution layer resulting in an output dimension of . For attribute-oriented description generation, we employ a 12-layer Multilingual BERT Kenton and Toutanova (2019) as the text encoder. After experimenting with various MLLMs, including mini-GPT4 (Zhu et al., 2023a) and Llava (Liu et al., 2023a), we ultimately selected Qwen-VL (Bai et al., 2023) for attribute-oriented description generation. For a fair comparison with other methods, we resized all images to as input resolution in both training and test phases throughout all experiments. The optimization of AttPL is done by AdamW Kingma and Ba (2014) with a learning rate of , weight decay , and the batch size 8 for maximally 20 epochs. All experiments are performed on Nvidia RTX-3090Ti GPU and our model is implemented in PyTorch 333http://xxx.com.
Metrics. To comprehensively evaluate performance, we adhere to the methodology of prior studies Wang et al. (2016); Ge et al. (2018); Chen et al. (2019) and present metrics including average per-class precision (CP), recall (CR), F1 (CF1), and the average overall precision (OP), recall (OR), F1 (OF1). Additionally, we include results for top-1, top-3, and top-5 tags. Higher values across all metrics indicate superior performance. Tags for each sticker image are deemed positive if their predicted probabilities surpass 0.5.
5.2 Compared Methods
We adopt comprehensive mainstream models for multi-tag recognition, building upon our proposed StickerTAG dataset: (1) Conventional methods: CNN Krizhevsky et al. (2012), ResNet-101 (He et al., 2016), ADD-GCN (Ye et al., 2020), ASL (Ridnik et al., 2021). (2) Transformer-based methods: Swin Liu et al. (2021b), ViT Dosovitskiy et al. (2020), BeiT Bao et al. (2021), DeiT Touvron et al. (2021a), CaiT Touvron et al. (2021b), CSRA Zhu and Wu (2021), Q2L Liu et al. (2021a), TSFormer Zhu et al. (2022), C-Tran Lanchantin et al. (2021). (3) Multi-modal large language methods: Mini-GPT4 Zhu et al. (2023b), LLaVA Liu et al. (2023b), Qwen-VL Bai et al. (2023).
@@ 5.3 Main Results
We examine the performance of our model and baselines in terms of each evaluation metric, as shown in Table 1. The experimental results indicate that the AttPL approach largely achieves the best results. Moreover, the significance tests of our AttPL over the baseline models demonstrate the effectiveness of our proposed method, presenting a statistically significant improvement based on most evaluation metrics with -value < 0.05. We also observe that the performance of Top-1 is better than Top-5, and overall it is inferior to individual categories. This indicates that in a multi-tag scenario, the model finds it challenging to accurately select all tags correctly, highlighting the inherent difficulty of this task.
Analysis of Transformer-based methods. ViT, DeiT, Cait, BeiT, and Swin models are based on the Transformer architecture, which is known for its self-attention mechanism, allowing the model to focus on different parts of the input sequence and weigh their importance concerning the current context. CSRA, as the strongest baseline model among the comparison models, benefits from generating class-specific features for each category. This is achieved by proposing a straightforward spatial attention score, which is then combined with class-agnostic average pooling features. This approach enables the model to more effectively capture the spatial distribution of objects from various categories within the image. However, these approaches are not well-suited for the multi-tag sticker recognition task due to their lack of consideration for sticker-specific characteristics, such as fine-grained visual and semantic features.
Analysis of multi-modal large language methods. We further analyze that the results of multi-modal large language models perform worse. On one hand, MLLMs have certain limitations in handling long texts, especially when dealing with a multitude of tags. When input into a MLLM for selection, it may not fully comprehend the context well. On the other hand, our observations from the output indicate that MLLMs tend to focus attention on the initial tags, potentially overlooking the subsequent ones. Thus, MLLMs exhibit suboptimal performance in multi-label scenarios. In the future, exploring ways to leverage MLLMs to enhance the performance of multi-tag or extreme multi-label recognition tasks is a worthwhile pursuit. Strategies such as employing larger models for data augmentation or incorporating reasoning explanations could be considered.
@@ 5.4 Ablation Study
To understand the influence of each component of our framework, we further conduct an ablation study. As shown in Table 1, the performances of all ablation models are worse than that of AttPL under all metrics, which demonstrates the necessity of each component in AttPL. The most significant decrease in performance occurs when excluding LoR, highlighting its paramount importance in our method. LoR plays a crucial role in guiding the model to focus more on the crucial information within stickers. Despite the performance drop without LoR, the model still outperforms other baselines, indicating that both prompt and penalty contribute to enhancing model performance to a certain extent. Furthermore, the thoughtful design of the prompt, incorporating sticker-specific features, contributes to the robustness of our model. This suggests that attributes of stickers play a pivotal role in the effective tag recognition of stickers.
Methods | CR | CF1 | OR | OF1 |
ours | 71.85 | 66.70 | 38.72 | 28.87 |
@@
Spatial | 69.44 | 62.88 | 37.87 | 26.51 |
Channel | 68.66 | 62.15 | 35.33 | 24.76 |
Self | 68.30 | 62.08 | 35.65 | 24.92 |
[…]
CaiT | 69.97 | 63.31 | 39.04 | 27.40 |
BeiT | 71.32 | 65.54 | 43.23 | 31.56 |
DeiT | 72.81 | 65.60 | 44.38 | 32.04 |
ViT | 73.31 | 65.81 | 45.08 | 32.42 |
@@ 5.5 Effect of Local Re-attention Module
On the other hand, we also explored various image encoders in LoR. It can be observed that our model achieves better performance when using ViT. This might be due to the ViT’s ability to capture global contextual information through self-attention in each image patch’s embedding, enhancing the model’s overall global perception when processing the entire image. Additionally, the LoR module designed in this paper is configured to deepen its focus on specific features and regions related to stickers, contributing to improved results. Furthermore, despite CaiT showing the least favorable performance as the image encoder, when considering this observation alongside Table 1, it becomes evident that our approach outperforms other baseline models.
@@ 5.6 Case Study
Several cases with ground truth tags and predicted tags inferred by our AttPL approach are depicted in Figure 6. Upon observation of examples (a) and (c), it becomes evident that sticker features are predominantly manifested in actions and expressions, establishing a direct correlation with the associated tags. For instance, a crying expression could be associated with tags like "I’m too distressed" or "Feeling wronged", while the act of holding one’s head and a contemplative facial expression may indicate tags such as "Thinking" or "I’ll think about it". The proposed method efficiently guides the model’s attention toward these crucial features, thereby improving recognition performance.
However, we also observed instances of incomplete recognition in example (b) of Figure 6 with our AttPL model. While it successfully infers the sticker as the "Autistic" tag, it struggles to recognize other tags in more depth. This challenge primarily arises from learning the similarities between different tags in multi-tag sticker recognition. Furthermore, for similar actions like the one representing the question "Why" in example (d), the model may mistakenly recognize it as the tag "I’ll think about it". This highlights the difficulty in distinguishing between similar tags and expressions, which is inherent in this task.

### v3

@@ Abstract
In real-world conversations, the diversity and ambiguity of stickers often lead to varied interpretations based on the context, necessitating the requirement for comprehensively understanding stickers and supporting multi-tagging. To address this challenge, we introduce StickerTAG, the first multi-tag sticker dataset comprising a collected tag set with 461 tags and 13,571 sticker-tag pairs111We believe that the release of this dataset will provide exciting research opportunities and encourage further research in sticker analysis. , designed to provide a deeper understanding of stickers. Recognizing multiple tags for stickers becomes particularly challenging due to sticker tags usually are fine-grained attribute aware. Hence, we propose an Attentive Attribute-oriented Prompt Learning method, i.e., Att2PL, to capture informative features of stickers in a fine-grained manner to better differentiate tags. Specifically, we first apply an Attribute-oriented Description Generation (ADG) module to obtain the description for stickers from four attributes. Then, a Local Re-attention (LoR) module is designed to perceive the importance of local information. Finally, we use prompt learning to guide the recognition process and adopt confidence penalty optimization to penalize the confident output distribution. Extensive experiments show that our method achieves encouraging results for all commonly used metrics.
@@
Bingbing Wang, Bin Liang, Chun-Mei Feng∗, Wangmeng Zuo, Zhixin Bai, Shijue Huang, Kam-Fai Wong, Xi Zeng, Ruifeng Xu∗
@@ 1 Introduction
To tackle this challenge, we design an Attentive Attribute-oriented Prompt Learning method, abbreviated as Att2PL, which begins with an Attribute-oriented Description Generation (ADG) module powered by a Multi-modal Large Language Model (MLLM) to derive the description of stickers based on the content, style, role, and action. Furthermore, the sticker is fed into a Local Re-attention (LoR) module for deriving patch-attentive embedding. Then, the recognition process is guided by prompt classification, where we initialize the soft prompt with attribute-oriented descriptions. Finally, we implement a confidence penalty optimization for precise and reliable recognition. We employ commonly used metrics to evaluate the performance of our model, and extensive experiments conducted on our StickerTAG dataset show that our approach significantly outperforms the strong baselines. Our contributions are summarized as follows:
@@
Experiments conducted on the StickerTAG dataset show that our model outperforms all baselines, clearly confirming its effectiveness.
@@ 2.2 Sticker-based Method
Unlike image-text pairs in vision-language datasets, stickers demand finer-grained emotion recognition and subject understanding due to their increased diversity and domain specificity. This has prompted the development of specialized approaches tailored to the unique attributes of stickers, as relying on generic vision-language models (VLMs) may prove suboptimal. CSMSA (Ge et al., 2022) introduces a Sticker-Aware Multi-modal Sentiment Analysis Model to tackle challenges in multimodality, inter-series variations, and multi-modal sentiment fusion. Zhao et al. (2023) delved into aligning visual and textual features with emotional cues and artificial painting. It investigates the feasibility of seamlessly integrating additional tools into Large Language Models (LLMs). However, simple classification methods are insufficient to address the challenges posed by multi-tag sticker recognition. This prompts us to delve into more fine-grained features of stickers and enhances the ability of the model to differentiate between tags more effectively.
@@ 3 StickerTAG Dataset
In this section, we provide a detailed description of our dataset construction process. Firstly, we discuss the process of tag construction, illustrated in Section 3.1. Next, we outline the procedure for collecting and annotating stickers in Section 3.2. Finally, we analyze the characteristics of our dataset in Section 3.3.
@@ 3.3 Characteristics
The final dataset contains 461 tags and 13,571 sticker-tag pairs annotated data, and the large-scale dataset is for future work. This dataset is the first annotated sticker dataset for multi-tag recognition to the best of our knowledge. We present the word cloud distribution of sticker tags in Figure 2 (a). We can conclude that certain tags prominently feature emotionally charged and conversational words, reflecting a diverse range of topics. Moreover, Figure 2 (b) shows the number samples per tag. It can be seen that our dataset exhibits a long-tailed distribution like most of the multi-label datasets. We also conduct studies on the tags of each sticker. According to our statistics, nearly 0.35% stickers have 6 tags, 2.67% of 5 tags, 9.23% of 4 tags, 14.99 % of 3 tags, 25.31% of 2 tags, and 47.45% of 1 tag. And the average tag length is 3.10 words.
@@ 4 Method
This section begins with a concise introduction to the multi-tag sticker recognition task, followed by a detailed explanation of our proposed Att2PL framework. As demonstrated in Figure 3, Att2PL consists primarily of four components: 1) Attribute-oriented Description Generation designs prompts based on the attributes of the stickers for MLLM to generate attribute-oriented descriptions. 2) Local Re-attention Module utilizes masked images in a novel manner to derive renewed attention for deriving patching-attentive embedding. 3) Prompt-based Classification exploits attribute-oriented descriptions to initialize the soft prompts and combines patch-attentive embedding to achieve the multi-tag prediction. 4) Confidence Penalty Optimization penalizes confident output distributions for implicit feedback of renewed attention.
@@ 4.5 Confidence Penalty Optimization
We deploy a confidence penalty optimization method to enhance the ability of LoR to capture important local information by penalizing confident output distributions. Specifically, we first obtain the original image representation by passing the image patch embedding through an image encoder. Then, concatenated with soft prompts initialized by four description representations to derive the original image representation , which is fed into the classifier, yielding the original confidence distribution. We compare the probability changes in reconstructed confidence and original confidence for the correct tag. We calculate the sum of the magnitudes of the decreases in reconstructed confidence relative to original confidence and incorporate it as a penalty into the loss.
@@
5 Experiment
5.1 Experimental Set-up
Implement details. For data participation, we divide our StickerTAG dataset into a training set, a validation set, and a testing set with a ratio of 8:1:1. Att2PL employs Swin Transformer Vaswani et al. (2017) as the image encoder, we apply a convolution layer resulting in an output dimension of . For attribute-oriented description generation, we employ a 12-layer Multilingual BERT Kenton and Toutanova (2019) as the text encoder. We utilize Qwen-VL (Bai et al., 2023) for attribute-oriented description generation. For a fair comparison with other methods, we resized all images to as input resolution in both training and test phases throughout all experiments. The optimization of Att2PL is done by AdamW Kingma and Ba (2014) with a learning rate of , weight decay , and the batch size 8 for maximally 20 epochs. All experiments are performed on Nvidia RTX-3090Ti GPU and our model is implemented in PyTorch 333http://xxx.com.
Metrics. To comprehensively evaluate performance, we adhere to the methodology of prior studies Wang et al. (2016); Ge et al. (2018); Chen et al. (2019) and present metrics including average per-class recall (CR), F1 (CF1), and the average overall recall (OR), F1 (OF1). Additionally, we include results for top-1, top-3, and top-5 tags. Higher values across all metrics indicate superior performance. Tags for each sticker image are deemed positive if their predicted probabilities surpass 0.5.
5.2 Compared Methods
We adopt comprehensive mainstream models for multi-tag recognition, building upon our proposed StickerTAG dataset: (1) Conventional methods: CNN Krizhevsky et al. (2012), ResNet-101 (He et al., 2016), ADD-GCN (Ye et al., 2020), ASL (Ridnik et al., 2021). (2) Transformer-based methods: Swin Liu et al. (2021b), ViT Dosovitskiy et al. (2020), BeiT Bao et al. (2021), DeiT Touvron et al. (2021a), CaiT Touvron et al. (2021b), CSRA Zhu and Wu (2021), Q2L Liu et al. (2021a), TSFormer Zhu et al. (2022), C-Tran Lanchantin et al. (2021), StickerCLIP Zhao et al. (2023), and LORA Liu et al. (2022). (3) Multi-modal large language methods: Mini-GPT4 Zhu et al. (2023b), LLaVA Liu et al. (2023b), Qwen-VL Bai et al. (2023). For MLLMs, we list the tag set, present the sticker image, and prompt them with " Please select the most suitable tags for the above sticker."
@@ 5.3 Main Results
We examine the performance of our model and baselines in terms of each evaluation metric, as shown in Table 1. The experimental results indicate that the Att2PL approach largely achieves the best results. Moreover, the significance tests of our Att2PL over the baseline models demonstrate the effectiveness of our proposed method, presenting a statistically significant improvement based on most evaluation metrics with -value < 0.05. We also observe that the performance of Top-1 is better than Top-5, and overall it is inferior to individual categories. This indicates that in a multi-tag scenario, the model finds it challenging to accurately select all tags correctly, highlighting the inherent difficulty of this task.
Analysis of Transformer-based methods. Transformer models are based on the Transformer architecture, which is known for its self-attention mechanism, allowing the model to focus on different parts of the input sequence and weigh their importance concerning the current context. CSRA, the strongest baseline, generates class-specific features using a spatial attention score combined with class-agnostic average pooling, effectively capturing spatial distributions. However, these approaches are not well-suited for the multi-tag sticker recognition task due to their lack of consideration for sticker-specific characteristics, such as fine-grained visual and semantic features.
Analysis of multi-modal large language methods. We find that multi-modal large language models (MLLMs) perform suboptimally, particularly with long texts and numerous tags. MLLMs often struggle to fully grasp context and tend to prioritize initial tags, overlooking subsequent ones. This results in poor performance in multi-label tasks. In the future, exploring ways to leverage MLLMs to enhance the performance of multi-tag or extreme multi-label recognition tasks is a worthwhile pursuit. Strategies such as employing larger models for data augmentation or incorporating reasoning explanations could be considered.
Experiments on SER30K Datasets. We also apply our Att2PL method on the existing SER30K dataset, as shown in Table 2. LORA Liu et al. (2022) captures local information in stickers through the attention mechanism. In contrast, our method not only focuses on local information but also explores the characteristics of stickers from different perspectives, delving deeper into their features to enhance classification performance.
@@ 5.4 Ablation Study
To understand the influence of each component of our framework, we further conduct an ablation study. As shown in Table 1, the performances of all ablation models are worse than that of Att2PL under all metrics, which demonstrates the necessity of each component in Att2PL. The most significant decrease in performance occurs when excluding LoR, highlighting its paramount importance in our method. LoR plays a crucial role in guiding the model to focus more on the crucial information within stickers. Despite the performance drop without LoR, the model still outperforms other baselines, indicating that both prompt and penalty contribute to enhancing model performance to a certain extent. Furthermore, the thoughtful design of the prompt, incorporating sticker-specific features, contributes to the robustness of our model. This suggests that attributes of stickers play a pivotal role in the effective tag recognition of stickers.
Methods | Accuracy |
TFN Zadeh et al. (2017) | 54.19 |
MCB Gao et al. (2016) | 58.18 |
PDANet-T Zhao et al. (2019) | 68.93 |
WSCNet-T Yang et al. (2018) | 69.45 |
LORA Liu et al. (2022) | 70.73 |
@@ 5.5 Effect of Local Re-attention Module
We also explored various image encoders in LoR. It can be observed that our model achieves better performance when using ViT. This might be due to the ViT’s ability to capture global contextual information through self-attention in each image patch’s embedding, enhancing the model’s overall global perception when processing the entire image. Additionally, the LoR module designed in this paper is configured to deepen its focus on specific features and regions related to stickers, contributing to improved results. Furthermore, despite CaiT showing the least favorable performance as the image encoder, when considering this observation alongside Table 1, it becomes evident that our approach outperforms other baseline models.
Methods | Top-1 | Top-3 | Top-5 | |||||||||
CR | CF1 | OR | OF1 | CR | CF1 | OR | OF1 | CR | CF1 | OR | OF1 | |
Ours | 71.85 | 66.70 | 38.72 | 28.87 | 68.69 | 70.14 | 32.50 | 48.21 | 70.17 | 68.39 | 37.19 | 36.60 |
[…]
Spatial | 69.44 | 62.88 | 37.87 | 26.51 | 66.98 | 68.45 | 31.20 | 47.01 | 68.13 | 66.43 | 35.38 | 34.78 |
Channel | 68.66 | 62.15 | 35.33 | 24.76 | 65.64 | 67.58 | 30.64 | 46.13 | 67.17 | 65.84 | 35.03 | 33.64 |
Self | 68.30 | 62.08 | 35.65 | 24.92 | 65.39 | 66.76 | 30.13 | 45.38 | 67.52 | 65.39 | 34.95 | 33.51 |
[…]
CaiT | 69.97 | 63.31 | 39.04 | 27.40 | 67.03 | 68.64 | 31.25 | 47.13 | 68.34 | 66.59 | 35.99 | 35.01 |
BeiT | 71.32 | 65.54 | 43.23 | 31.56 | 67.81 | 69.42 | 32.08 | 47.76 | 69.79 | 67.85 | 37.08 | 35.97 |
DeiT | 72.81 | 65.60 | 44.38 | 32.04 | 68.99 | 70.07 | 32.34 | 48.25 | 69.94 | 68.54 | 37.03 | 36.64 |
ViT | 73.31 | 65.81 | 45.08 | 32.42 | 69.32 | 70.54 | 32.58 | 48.34 | 70.21 | 68.75 | 37.47 | 37.02 |
@@ 5.6 Case Study
Several cases with ground truth tags and predicted tags inferred by our Att2PL approach are depicted in Figure 5. Upon observation of examples (a) and (c), it becomes evident that sticker features are predominantly manifested in actions and expressions, establishing a direct correlation with the associated tags. For instance, a crying expression could be associated with tags like "I’m too distressed" or "Feeling wronged", while the act of holding one’s head and a contemplative facial expression may indicate tags such as "Thinking" or "I’ll think about it". The proposed method efficiently guides the model’s attention toward these crucial features, thereby improving recognition performance.
Error analysis. We also observed instances of incomplete recognition in example (b) of Figure 5 with our Att2PL model. While it successfully infers the sticker as the "Autistic" tag, it struggles to recognize other tags in more depth. This challenge primarily arises from learning the similarities between different tags in multi-tag sticker recognition. Furthermore, for similar actions like the one representing the question "Why" in example (d), the model may mistakenly recognize it as the tag "I’ll think about it". This highlights the difficulty in distinguishing between similar tags and expressions, which is inherent in this task.